OpenStudy (anonymous):

Karen, her brother Dan, Dana, and her sister Randi own a pizza restaurant. Karen and Dan each own 2/7 of the restaurant. The remaining share is owned equally by Dana and Randi. Exactly what fraction of the business does Dana own?

OpenStudy (anonymous):

3/14 ths

OpenStudy (anonymous):

7/7 - 2*(2/7) = 7/7 - 4/7 = 3/7 Can you find half of 3/7?